OpenStudy (anonymous):

The line containing the longer diagonal of a quadrilateral whose vertices are A (2, 2), B(-2, -2), C(1, -1), and D(6, 4).

OpenStudy (anonymous):

calculate AC and BD using distance formula and the longer one will be the line containing the longer diagonal of the quadrilateral

OpenStudy (anonymous):

so x-y=9?

OpenStudy (anonymous):

or 2x+1y=9

OpenStudy (anonymous):

AC = sqrt[(x2 - x1)^2 + (y2 - y1)^2]

OpenStudy (anonymous):

AC = sq.rt[(1 - 2)^2 + (-1 -2)^2] ac = sqrt[(-1)^2 + (-3)^2] AC = sqrt[1 + 9] ac = sqrt(10)

OpenStudy (anonymous):

similarly calculate BD

OpenStudy (anonymous):

B(-2, -2), D(6, 4). BD = sq.rt{[6 - (-2)]^2 + [4 -(-2)]^2} BD = sqrt[(6 + 2)^2 + (4 + 2)^2] BD = sqrt[(8)^2 + (6)^2] BD = sqrt(64 + 36) BD = sqrt(100) BD = 10 so BD is the longer one

OpenStudy (anonymous):

can you help me with another equation

OpenStudy (anonymous):

sure, go ahead..

OpenStudy (anonymous):

Indicate the equation of the given line in standard form. The line containing the median of the trapezoid whose vertices are R(-1, 5) , S(l, 8), T(7, -2), and U(2, 0). This is the question... and for the answer i got was: 13+x-y=48? can you tell me where i went wrong

OpenStudy (anonymous):

We can first find the midpoints of the legs RU and ST using the Midpoint formula of a line segment. Then we find the distance between them by using the formula for Distance between Two Points.

OpenStudy (anonymous):